1/2 log x–0.25 =3/4
Answer by stanbon(75887)

1/2 log x–0.25 =3/4
----------------------
(1/2)log(x-0.25) = 3/4
----------------------
log(x-0.25) = sqrt(3/4)
-----
x-0.25 = 10^(sqrt(3/4))
-----
x - 0.25 = 7.35
-----
x = 7.6
===============
Cheers,
Stan H.
---------------